OpenStudy (anonymous):

I just want to make sure that I am indeed correct! ~Slope~ Which of the following is the slope between the two points (5, 1) and (4, 6)? 1/5 -1/5 5 -5 I believe it is 5!

OpenStudy (luigi0210):

Well let's plug it all in and see: \[\huge \frac{y_{2}-y_{1}}{x_{2}-x_{1}}=\frac{6-1}{4-5}=-\frac{5}{1}\] You were so close Lacey :3

OpenStudy (anonymous):

Thanks Lui :P

OpenStudy (anonymous):

I can always turn to you :)

OpenStudy (luigi0210):

Anytime, now if you would excuse me, I have some laptops to look at.

OpenStudy (anonymous):

oh fine! jk Did I tell you I signed up to be an ambie...

OpenStudy (luigi0210):

Hmm, actually no you didn't.

OpenStudy (anonymous):

I did the form thingy :P

OpenStudy (luigi0210):

Oh, awesome: \[\huge \mathbb {\color{red}{I}~\color{orange}{WISH}~\color{yellow}{YOU}~\color{navy}{THE}~\color{blue}{BEST}~\color{aqua}{OF}~\color{green}{LUCK}} {\ddot\smile}\]

OpenStudy (luigi0210):

\[\Huge \color{red}{\heartsuit}\]